OpenStudy (pokemon23):

simplify: (-5xy^2)-(-3xy^2) simplify: (-5xy^2)-(-3xy^2) @Mathematics

OpenStudy (anonymous):

here

OpenStudy (anonymous):

combine like terms by computing −5−(−3)=−5+3=−2

OpenStudy (anonymous):

combine like terms by computing −5−(−3)=−5+3=−2

OpenStudy (anonymous):

Did it help a little_ I hope Good luck!

OpenStudy (pokemon23):

-2xy^4?

OpenStudy (anonymous):

Yep i think

OpenStudy (pokemon23):

lol